1)     Consider a consumer with a utility function U (x, y) =X2 + Y2, the consumer intends to spend birr 80 on the two goods and price of good X and price of good Y are birr 2 and birr 4, respectively

A.   Calculate the optimum consumers consumption amount of X and Y

B.    Find the maximum utility that consumer obtain from consuming the two goods?

C.     Calculate MRSxy at equilibrium, and interpret your result


1
Expert's answer
2022-01-19T10:18:50-0500

U(X,Y)= "X^2+Y^2"

m=80

"P_X=2"

"P_y=4"

"\frac{MU_X}{MU_Y}=\frac{P_x}{P_y}"

"MU_x= 2X"

"MU_Y= 2Y"

"\frac{2X}{2Y}= \frac{2}{4}\implies y=2x"


From the Budget constraint

"2X^2+4Y^2=80"


"2X^2+4(2X)^2=80"

"X^2=4.4\implies X=2.1"

"Y= (2.1)^2= 4.4"


Combined Utility Maximization

"U(X,Y)= X^2+Y^2"

"= (2.1)^2+ (4.4)^2= 23.76"


"MRS=\frac{\delta U}{(\delta_x\delta_y)}"

"MRS=\frac{4}{(x+y)^2}=\frac{4}{(2.1+4.4)^2}"

"= 0.0946"


There is a very low substitution effect